# PolyhedronNet: Representation Learning for Polyhedra with Surface-attributed Graph

### v1

@@ 1 Introduction
In mathematics and computational geometry, a polyhedron is defined as a three-dimensional (3D) solid formed by flat polygon faces joined at edges and vertices. Ubiquitous geometric shapes can be precisely and efficiently modeled as polyhedra, ranging from basic 3D shapes (e.g., cubic, pyramid, and truncated tetrahedron) to compositions of them (e.g., shapes of buildings, furniture, and digital objects in CAD) as exemplified Figure 1 (a). In the real world, there are many tasks surrounding polyhedra such as classification (e.g., convex or concave); clustering polyhedra into different types (e.g., Platonic solids and prisms ); as well as generation and optimization (e.g., use faceted facades to break up flat surfaces) of polyhedra for design needs. However, the raw form of polyhedra cannot be directly input into machine learning models which require structured formats such as vectors, tensors, etc. Hence, a fundamental upstream task is to map a polyhedron into a vector representation, namely polyhedra representation learning, which is the focus of this paper.
@@ 2.2 Polyhedral Representation Learning
Recent advancements in the field of polyhedral geometry representation learning have been significant. Traditional feature engineering approaches (Pham et al., 2010; Yan et al., 2019; He et al., 2018) transform polygonal shapes into predefined shape descriptors. GNNs are utilized to improve handling of spatial relationships and structural complexities (Qi et al., 2017b; Shi & Rajkumar, 2020; Wang et al., 2019). However, these descriptors tend to oversimplify the data, failing to capture the complete spectrum of shape information and requiring substantial domain expertise for their creation. They struggle with the variability and complexity of polygonal shapes, which limits their generalizability. Polygon shape encoding methods (van’t Veer et al., 2019; Mai et al., 2023; Yan et al., 2021), have demonstrated their effectiveness in shape classification and retrieval tasks. While beneficial for certain types of analysis, these methods do not fully meet the needs of polyhedral representation learning that requires capturing complex topological relationships between polygonal geometries. In relation to polyline representation learning (Jiang et al., 2021; 2022), these methods focus on processing continuous lines and curves that delineate the boundaries and configurations of shapes in spatial data. However, when dealing with intersecting or overlapping geometric structures, the handling of their topological relationships can be complex and may not be sufficient to accurately capture more intricate curves and nonlinear structures. Another category of research focuses on polyhedron generation. Gillsjö et al. (2023) extracts polygons from images by using heterogeneous graphs and wireframes to learn feature space. Zorzi & Fraundorfer (2023) utilizes edge-aware GNNs to enhance polygon detection accuracy and applicability in scene parsing by considering both node and edge features. Antonietti et al. (2024) enhance the analysis of geometric structures by maintaining mesh quality and improving computational processes, as demonstrated in multigrid solvers and scene parsing tasks.
@@ Appendix F Performance of more comparison methods and ablated models
We conduct additional experiments to evaluate our method against more comparison methods and ablated models. First, we compare different aggregation strategies, including mean and max aggregators, where results show that max aggregation generally achieves superior performance on MNIST-C and ShapeNet-P datasets, while mean aggregation performs better on Building and ModelNet-P datasets. To validate the effectiveness of our proposed heterogeneous geometric message passing mechanism, we conduct an ablation study (our w/o hetero) where we remove the heterogeneous message-passing modules. The significant performance drop demonstrates the importance of these components in capturing complex geometric relationships. We also compare our method with established graph learning methods, including HGT and HAN. The results show that our approach substantially outperforms these methods across all datasets, indicating the advantage of our design. Furthermore, we benchmark against recent state-of-the-art point cloud methods, including LocoTrans and RISurConv. The experimental results demonstrate that our method achieves superior performance on most datasets, particularly showing significant improvements on MNIST-C and Building datasets. This suggests that our approach better captures the inherent geometric structure of 3D shapes compared to point-based methods.

### v2

@@ 1 Introduction
In mathematics and computational geometry Yu et al. (2025), a polyhedron is defined as a three-dimensional (3D) solid formed by flat polygon faces joined at edges and vertices. Ubiquitous geometric shapes can be precisely and efficiently modeled as polyhedra, ranging from basic 3D shapes (e.g., cubic, pyramid, and truncated tetrahedron) to compositions of them (e.g., shapes of buildings, furniture, and digital objects in CAD) as exemplified Figure 1 (a). In the real world, there are many tasks surrounding polyhedra such as classification (e.g., convex or concave); clustering polyhedra into different types (e.g., Platonic solids and prisms ); as well as generation and optimization (e.g., use faceted facades to break up flat surfaces) of polyhedra for design needs. However, the raw form of polyhedra cannot be directly input into machine learning models which require structured formats such as vectors, tensors, etc. Hence, a fundamental upstream task is to map a polyhedron into a vector representation, namely polyhedra representation learning, which is the focus of this paper.
@@ 2.2 Polyhedral Representation Learning
Recent advancements in the field of polyhedral geometry representation learning have been significant. Traditional feature engineering approaches (Pham et al., 2010; Yan et al., 2019; He et al., 2018) transform polygonal shapes into predefined shape descriptors. GNNs are utilized to improve the handling of spatial relationships and structural complexities (Qi et al., 2017b; Shi & Rajkumar, 2020; Wang et al., 2019). However, these descriptors oversimplify the data, failing to capture the complete spectrum of shape information. Polygon shape encoding methods (van’t Veer et al., 2019; Mai et al., 2023; Yan et al., 2021), have demonstrated their effectiveness in shape classification and retrieval tasks. While beneficial for certain types of analysis, these methods do not fully meet the needs of polyhedral representation learning that requires capturing complex topological relationships between polygonal geometries. PolygonGNN Yu et al. (2024) is the first GNN-based method that captures both multi-polygon relationships and individual polygon shape information. However, it is designed specifically for 2D shapes. In relation to polyline representation learning (Jiang et al., 2021; 2022), current works focus on processing continuous lines and curves that delineate the boundaries and configurations of shapes in spatial data. Another category of research focuses on polyhedron generation. Gillsjö et al. (2023) extracts polygons from images by using heterogeneous graphs and wireframes to learn feature space. Zorzi & Fraundorfer (2023) utilizes edge-aware GNNs to enhance polygon detection accuracy and applicability in scene parsing by considering both node and edge features.
@@ 6 Conclusion
Acknowledgments
This work was partially supported by the NSF Grant No. 2007716, No. 2007976, No. 1942594, No. 1907805, Cisco Faculty Research Award, and NIH Grant No. R01AG089806.
@@ Appendix F Performance of more comparison methods and ablated models
We conduct additional experiments to evaluate our method against more comparison methods and ablated models. First, we compare different aggregation strategies, including mean and max aggregators, where results show that max aggregation generally achieves superior performance on MNIST-C and ShapeNet-P datasets, while mean aggregation performs better on Building and ModelNet-P datasets. To validate the effectiveness of our proposed heterogeneous geometric message passing mechanism, we conduct an ablation study (our w/o hetero) where we remove the heterogeneous message-passing modules. The significant performance drop demonstrates the importance of these components in capturing complex geometric relationships. We also compare our method with established graph learning methods, including HGT and HAN. The results show that our approach substantially outperforms these methods across all datasets, indicating the advantage of our design. Furthermore, we benchmark against recent state-of-the-art point cloud methods, including LocoTrans Chen et al. (2024) and RISurConv Zhang et al. (2024). The experimental results demonstrate that our method achieves superior performance on most datasets, particularly showing significant improvements on MNIST-C and Building datasets. This suggests that our approach better captures the inherent geometric structure of 3D shapes compared to point-based methods.